OpenStudy (anonymous):

Find the slope of the tangent line to the graph of f at the given point. f(x) = x2 + 5x at (4, 36)

OpenStudy (anonymous):

\[f(x) = x^2 + 5x\\f(x)'=2x+5\] The slope is: \[f(4)'=2(4)+5=13\] I'm sure you will find the tangent line c:

OpenStudy (anonymous):

Thank You, now I understand.